# MoH: Multi-Head Attention as Mixture-of-Head Attention

### v2

@@ 4.2 DiT for Class-Conditional Image Generation
Model Settings.    For Diffusion models with Transformers (DiT) (Peebles & Xie, 2023), we only replace the standard multi-head attention with our MoH in MoH-DiT models, while keeping all other training parameters identical to DiT. We use the ImageNet-1K dataset for class-conditional image generation at a resolution of 256$\times$256. To evaluate generation performance, we use Frechet Inception Distance (FID(Heusel et al., 2017) to assess overall sample quality, Precision and Recall (Kynkäänniemi et al., 2019) to measure fidelity and diversity separately, and sFID (Nash et al., 2021) as a metric that better captures spatial relationships than FID. Moreover, we use Inception Score (IS(Salimans et al., 2016) as another metric for fidelity.
Training Details.    Following DiT, the final linear layer is initialized with zeros, and all other layers follow standard ViT weight initialization. We train all models using the AdamW optimizer (Loshchilov & Hutter, 2017) with a constant learning rate of 1e-4, no weight decay, and a batch size of 256, applying horizontal flips for data augmentation. Following DiT, we employ the Exponential Moving Average (EMA) of MoH-DiT weights during training with a decay rate of 0.9999, generating all images using the EMA model. We use an off-the-shelf pre-trained variational autoencoder (Kingma, 2013) model from Stable Diffusion (Rombach et al., 2022). Following TransNeXt, our attention-head activation budget is unevenly distributed across layers, with fewer attention heads activated in the shallow layers and more in the deeper layers.
@@ 4.3 Training LLMs from Scratch
Model Settings.    For training LLMs from scratch, we use Megatron (Shoeybi et al., 2019), an open-source training code, as the training framework. Please refer to the Appendix for detailed hyper-parameter settings (Tab. C) of various MoH-LLMs. The evaluation is performed on multiple benchmarks using the Eleuther AI Language Model Evaluation Harness (Gao et al., 2024), a unified framework for testing generative language models. Since the parameters are only about 0.2B for the smallest model, we select 6 simple benchmarks as the metric. Specifically, we report 0-shot accuracy on SciQ (Welbl et al., 2017), PIQA (Bisk et al., 2020), WinoGrande (Sakaguchi et al., 2021), OpenbookQA (Mihaylov et al., 2018), LogiQA (Liu et al., 2020), and TruthfulQA (Lin et al., 2022).
Training Details.    We only use public datasets for training, ensuring accessibility for academic research. Specifically, we sample from the RedPajama (Computer, 2023), Dolma (Soldaini et al., 2024), and Pile (Gao et al., 2020) datasets according to different sampling probabilities. Please refer to the Appendix for detailed sample ratios. Following previous works (Jin et al., 2025), we utilize the tokenizer from LLaMA2 (Touvron et al., 2023), which contains 65,536 vocabulary tokens.
@@ 4.4 Continue-Tuning LLaMA3-8B
Model Settings.    To significantly enhance the applicability of the proposed MoH method, we also attempt to further continue-tune pre-trained multi-head attention models, such as LLaMA3-8B, into MoH models. However, this presents three challenges. (i) Determining the shared attention heads: We simply select the first 16 attention heads of each layer as shared heads. (ii) Adding head routers: Integrating a randomly initialized router into the pre-trained model without compromising its original performance requires careful training techniques. To address this, we propose a parameter-free router that determines routing scores using the $\ell_{2}$ norm of the query of each attention head. (iii) Weighting attention heads: We observe that weighting the attention head outputs significantly alters the distribution of the output of the attention layer, which necessitates a large amount of training data to restore the original performance. To tackle this, we quantize the routing score and use the straight-through estimator (Bengio et al., 2013; Liu et al., 2022) to back-propagate the gradients through the sparsity function. Specifically, given the input token $\bm{x}$, we employ a quantizer for activation routing scores, with its forward pass formulated as:
@@
where $g_{i}$ denotes the real-valued routing score. This approximation function significantly mitigates the issue of gradient vanishing (Wang et al., 2024). Similar to training LLMs from scratch, we also use Megatron (Shoeybi et al., 2019), an open-source training code, as the training framework.
@@
Training Details.    We find that if there is a discrepancy between the continue-training data and the original training data distribution of the model, the performance of the model may fluctuate wildly at the beginning of the training process. Since we are unable to have access to the raw training data of LLaMA3, we address these potential performance fluctuations by dividing the training process into two stages. In the first stage, we continue-tune the original LLaMA3-8B model using 300B tokens to adapt the model to our dataset. In the second stage, we continue-tune this adapted model into our proposed MoH model with 100B tokens. We utilize the lm-evaluation-harness package to assess performance on a comprehensive suite of downstream tasks: (i) Following Pythia (Biderman et al., 2023), we report 0-shot accuracy on LAMBADA (Paperno et al., 2016), LogiQA (Liu et al., 2020), PIQA (Bisk et al., 2020), SciQ (Welbl et al., 2017), and WinoGrande (Sakaguchi et al., 2021). (ii) We report the accuracy of Chinese tasks, including 5-shot CEVAL (Huang et al., 2023) and 5-shot CMMLU (Li et al., 2023a). (iii) We report the accuracy of tasks from the Open LLM Leaderboard (Beeching et al., 2023), including 10-shot HellaSwag (Zellers et al., 2019), 25-shot ARC Challenge (ARC-C) (Clark et al., 2018), and 5-shot MMLU (Hendrycks et al., 2021). (iv) We report the exact match score for 32-shot Natural Questions (NQ) (Kwiatkowski et al., 2019) and the accuracy for 32-shot BoolQ (Clark et al., 2019). (v) We report the exact match score for 8-shot GSM8K (Cobbe et al., 2021) to evaluate the math ability. (vi) Moreover, we report 0-shot accuracy on TruthfulQA (Lin et al., 2022) to assess the ability to generate truthful answers.
@@ 4.5 Ablative Analysis
Effect of the Shared Heads Ratio among Activated Heads.    In Tab. 7, we provide the ablation study on the shared heads ratio among activated heads. We find that model performance remains relatively consistent across a wide range of shared heads ratios (from 13.9% to 74.0%). These results indicate that the performance of the model is stable as long as the shared heads ratio is not extreme. From another perspective, shared heads can be viewed as a form of Soft MoE (Puigcerver et al., 2024). Based on the findings from the Soft MoE paper (Puigcerver et al., 2024), we recommend using a higher ratio of shared heads among the activated heads (greater than 40%).
@@ 5 Discussion
The Difference between MoH and MoA.    We clarify the differences between MoH and MoA (Zhang et al., 2022) from the following three aspects. First, in terms of motivation, the goal of MoH is to improve the efficiency and performance of the attention mechanism without increasing the number of parameters. In contrast, MoA shares the motivation of MoE, which is to expand model parameters while keeping inference costs low. Therefore, the model settings of MoH are more stringent than those of MoA. Second, in terms of methodology, our MoH introduces shared heads and two-stage routing to enhance the standard MoE method. More importantly, we show that pre-trained multi-head attention models can be further continue-tuned into our MoH models, greatly improving the applicability of the proposed MoH method. In contrast, MoA directly combines multi-head attention with MoE. Due to the adoption of shared keys and values, MoA must be trained from scratch, which limits its applicability. Finally, in terms of model frameworks, our MoH is validated across various popular model frameworks and tasks, including ViT, DiT, and decoder-only LLMs, while MoA is only validated for language tasks.
@@ 6 Conclusion
In this work, we introduce MoH, a promising alternative to multi-head attention. MoH enables each token to adaptively select the appropriate attention heads, improving both model performance and inference efficiency without increasing the number of parameters. Extensive experiments across various popular model frameworks, including ViT, DiT, and LLMs, demonstrate that MoH outperforms multi-head attention, even when using only 50$\sim$90of the attention heads. This work represents a promising step toward advanced and efficient attention-based models, which may be helpful to both the research and industrial communities.
@@ A.1 Why is MoH Superior to Vanilla Multi-Head Attention?
In contrast, in MoH, routed heads are trained only on smaller subsets of data specifically assigned to them. In MoH’s routing mechanism, the data is divided into $h-h_{s}$ subsets $\{D_{1},D_{2},...,D_{h-h_{s}}\}$, with each subset corresponding to a routed head. Besides, the routing score for each attention head acts as an adaptive adjustment to the learning rate, enabling the attention heads in MoH to specialize more effectively. Given a minibatch of data $D$ and the router $G(*)$, the gradient of each routed head in MoH can be written as $\mathbb{E}_{x\in D_{i}}[G(x)_{i}\frac{\partial\mathcal{L}(x)}{\partial h_{i}}]$. The gradient of each shared head in MoH can be written as $\mathbb{E}_{x\in D}[G(x)_{i}\frac{\partial\mathcal{L}(x)}{\partial h_{i}}]$. As shown in Tab. A, the routing mechanism and adaptive weights in MoH enable attention heads to specialize more effectively compared to standard multi-head attention.
@@
Table B: The similarity of attention patterns and output features among attention heads. Given a pair of attention score matrices $A$ and $A^{\prime}$, we calculate the similarity of attention patterns as $1-\frac{1}{2}\mathbb{E}[||A-A^{\prime}||_{1}]$. Since attention scores form a probability distribution for each query, the similarity is always between 0 to 1.
@@ A.2 Limitations and Future Work
Lower Activation Rate.    Currently, MoH outperforms multi-head attention by utilizing only 50$\sim$90of the attention heads. However, this is still a relatively high proportion. Future work could aim to further optimize MoH, reducing head activation to less than 50%.
Multimodal Inputs.    Effectively processing information from multiple modalities in the attention mechanism remains an open question. Recent work (Wan et al., 2024) has shown that visual and textual tokens exhibit distinct attention patterns in multi-head attention. Future work could explore the attention patterns of MoH with different modal inputs, for example within multimodal large language models (Jin et al., 2024b; Lin et al., 2023, 2024; Liu et al., 2024; Jin et al., 2023, 2024a).
@@ B.1 ViT for Image Classification
Training Details.    Our MoH-ViT models are trained for 300 epochs using automatic mixed precision across 8 GPUs. We follow the training strategy of TransNeXt, which includes various data augmentation techniques, including Random Augmentation (Cubuk et al., 2020), Mixup (Zhang, 2017), CutMix (Yun et al., 2019), and Random Erasing (Zhong et al., 2020). We also apply Label Smoothing (Szegedy et al., 2016) and DropPath (Huang et al., 2016) to regularize our models. We optimize our models using AdamW optimizer (Loshchilov & Hutter, 2017) with a gradient clipping norm of 1.0 and a weight decay of 0.05. The initial learning rate is set to 1e-3, with a 5-epoch warm-up starting at 1e-6. A cosine learning rate scheduler (Loshchilov & Hutter, 2016) is employed to decay the learning rate. During training, images are randomly cropped to a size of 224$\times$224. It is worth noting that we do not use Exponential Moving Average (EMA) weights.
@@ B.2 DiT for Class-Conditional Image Generation
Training Details.    Following DiT, the final linear layer is initialized with zeros, and all other layers follow standard ViT weight initialization. We train all models using the AdamW optimizer (Loshchilov & Hutter, 2017) with a constant learning rate of 1e-4, no weight decay, and a batch size of 256, applying horizontal flips for data augmentation. Following DiT, we employ the Exponential Moving Average (EMA) of MoH-DiT weights during training with a decay rate of 0.9999, generating all images using the EMA model. We use an off-the-shelf pre-trained variational autoencoder (Kingma, 2013) model from Stable Diffusion (Rombach et al., 2022). Following TransNeXt, our attention-head activation budget is unevenly distributed across layers, with fewer attention heads activated in the shallow layers and more in the deeper layers.
@@ B.3 Training LLMs from Scratch
Model Settings.    For training LLMs from scratch, we use Megatron (Shoeybi et al., 2019), an open-source training code, as the training framework. The detailed hyper-parameter settings of various MoH-LLMs are shown in Tab. C.
@@
  • The RedPajama (Computer, 2023) includes training data from seven domains: CommonCrawl, C4, Github, Wikipedia, Books, ArXiv, and StackExchange.
  • The Dolma (Soldaini et al., 2024), a large and diverse open English text corpus, contains 3 trillion tokens sampled from seven sources, including web pages from Common Crawl, code from The Stack, curated web data from C4 (Raffel et al., 2020), social media conversations from Reddit, academic papers from PeS2o, public domain books from Project Gutenberg, and comprehensive content from Wikipedia and Wikibooks.
  • The Pile (Gao et al., 2020), an open-source English text corpus for training large language models, includes 22 diverse, publicly available datasets such as Wikipedia, NIH ExPorter, ArXiv, Books3, BookCorpus2, OpenSubtitles, YoutubeSubtitles, and Enron Emails.
@@
Training Hyper-Parameters.    Tab. E shows the detailed training hyper-parameters of MoH-LLMs. Specifically, all MoH-LLMs are trained with the AdamW optimizer (Loshchilov & Hutter, 2017), using a batch size of 4 million tokens with a sequence length of 2048. The final learning rate is set to 10% of the maximum. During training, a weight decay of 0.1 and gradient clipping of 1.0 are applied. For LLM-S and MoH-LLM-S, the maximum learning rate is set to 3e-4. For LLM-B and MoH-LLM-B, the maximum learning rate is set to 5e-4.

### v3

@@ 4.2 DiT for Class-Conditional Image Generation
Model Settings.   For Diffusion models with Transformers (DiT) (peebles2023scalable), we only replace the standard multi-head attention with our MoH in MoH-DiT models, while keeping all other training parameters identical to DiT. We use the ImageNet-1K dataset for class-conditional image generation at a resolution of $\times256$. To evaluate generation performance, we use Frechet Inception Distance (FID(heusel2017gans) to assess overall sample quality, Precision and Recall (kynkaanniemi2019improved) to measure fidelity and diversity separately, and sFID (nash2021generating) as a metric that better captures spatial relationships than FID. Moreover, we use Inception Score (IS(salimans2016improved) as another metric for fidelity.
Training Details.   Following DiT, the final linear layer is initialized with zeros, and all other layers follow standard ViT weight initialization. We train all models using the AdamW optimizer (loshchilov2017decoupled) with a constant learning rate of 1e-4, no weight decay, and a batch size of 256, applying horizontal flips for data augmentation. Following DiT, we employ the Exponential Moving Average (EMA) of MoH-DiT weights during training with a decay rate of 0.9999, generating all images using the EMA model. We use an off-the-shelf pre-trained variational autoencoder (kingma2013auto) model from Stable Diffusion (rombach2022high). Following TransNeXt, our attention-head activation budget is unevenly distributed across layers, with fewer attention heads activated in the shallow layers and more in the deeper layers.
@@ 4.3 Training LLMs from Scratch
Model Settings.   For training LLMs from scratch, we use Megatron (shoeybi2019megatron), an open-source training code, as the training framework. Please refer to the Appendix for detailed hyper-parameter settings (Tab. C) of various MoH-LLMs. The evaluation is performed on multiple benchmarks using the Eleuther AI Language Model Evaluation Harness (eval-harness), a unified framework for testing generative language models. Since the parameters are only about 0.2B for the smallest model, we select 6 simple benchmarks as the metric. Specifically, we report 0-shot accuracy on SciQ (welbl2017crowdsourcing), PIQA (bisk2020piqa), WinoGrande (sakaguchi2021winogrande), OpenbookQA (OpenBookQA2018), LogiQA (liu2020logiqa), and TruthfulQA (lin-etal-2022-truthfulqa).
Training Details.   We only use public datasets for training, ensuring accessibility for academic research. Specifically, we sample from the RedPajama (together2023redpajama), Dolma (dolma), and Pile (pile) datasets according to different sampling probabilities. Please refer to the Appendix for detailed sample ratios. Following previous works (jin2024moe++), we utilize the tokenizer from LLaMA2 (touvron2023llama2), which contains 65,536 vocabulary tokens.
@@ 4.4 Continue-Tuning LLaMA3-8B
Model Settings.   To significantly enhance the applicability of the proposed MoH method, we also attempt to further continue-tune pre-trained multi-head attention models, such as LLaMA3-8B, into MoH models. However, this presents three challenges. (i) Determining the shared attention heads: We simply select the first 16 attention heads of each layer as shared heads. (ii) Adding head routers: Integrating a randomly initialized router into the pre-trained model without compromising its original performance requires careful training techniques. To address this, we propose a parameter-free router that determines routing scores using the $\ell_{2}$ norm of the query of each attention head. (iii) Weighting attention heads: We observe that weighting the attention head outputs significantly alters the distribution of the output of the attention layer, which necessitates a large amount of training data to restore the original performance. To tackle this, we quantize the routing score and use the straight-through estimator (bengio2013estimating; liu2022nonuniform) to back-propagate the gradients through the sparsity function. Specifically, given the input token $\bm{x}$, we employ a quantizer for activation routing scores, with its forward pass formulated as:
@@
where $g_{i}$ denotes the real-valued routing score. This approximation function significantly mitigates the issue of gradient vanishing (wang2024q). Similar to training LLMs from scratch, we also use Megatron (shoeybi2019megatron), an open-source training code, as the training framework.
@@
Training Details.   We find that if there is a discrepancy between the continue-training data and the original training data distribution of the model, the performance of the model may fluctuate wildly at the beginning of the training process. Since we are unable to have access to the raw training data of LLaMA3, we address these potential performance fluctuations by dividing the training process into two stages. In the first stage, we continue-tune the original LLaMA3-8B model using 300B tokens to adapt the model to our dataset. In the second stage, we continue-tune this adapted model into our proposed MoH model with 100B tokens. We utilize the lm-evaluation-harness package to assess performance on a comprehensive suite of downstream tasks: (i) Following Pythia (biderman2023pythia), we report 0-shot accuracy on LAMBADA (paperno2016lambada), LogiQA (liu2020logiqa), PIQA (bisk2020piqa), SciQ (welbl2017crowdsourcing), and WinoGrande (sakaguchi2021winogrande). (ii) We report the accuracy of Chinese tasks, including 5-shot CEVAL (huang2024c) and 5-shot CMMLU (li2023cmmlu). (iii) We report the accuracy of tasks from the Open LLM Leaderboard (open-llm-leaderboard-v1), including 10-shot HellaSwag (zellers2019hellaswag), 25-shot ARC Challenge (ARC-C) (clark2018think), and 5-shot MMLU (hendrycks2020measuring). (iv) We report the exact match score for 32-shot Natural Questions (NQ) (kwiatkowski2019natural) and the accuracy for 32-shot BoolQ (clark2019boolq). (v) We report the exact match score for 8-shot GSM8K (cobbe2021training) to evaluate the math ability. (vi) Moreover, we report 0-shot accuracy on TruthfulQA (lin-etal-2022-truthfulqa) to assess the ability to generate truthful answers.
@@ 4.5 Ablative Analysis
Effect of the Shared Heads Ratio among Activated Heads.   In Tab. 7, we provide the ablation study on the shared heads ratio among activated heads. We find that model performance remains relatively consistent across a wide range of shared heads ratios (from 13.9% to 74.0%). These results indicate that the performance of the model is stable as long as the shared heads ratio is not extreme. From another perspective, shared heads can be viewed as a form of Soft MoE (puigcerver2023sparse). Based on the findings from the Soft MoE paper (puigcerver2023sparse), we recommend using a higher ratio of shared heads among the activated heads (greater than 40%).
@@ 5 Discussion
The Difference between MoH and MoA.   We clarify the differences between MoH and MoA (zhang2022mixture) from the following three aspects. First, in terms of motivation, the goal of MoH is to improve the efficiency and performance of the attention mechanism without increasing the number of parameters. In contrast, MoA shares the motivation of MoE, which is to expand model parameters while keeping inference costs low. Therefore, the model settings of MoH are more stringent than those of MoA. Second, in terms of methodology, our MoH introduces shared heads and two-stage routing to enhance the standard MoE method. More importantly, we show that pre-trained multi-head attention models can be further continue-tuned into our MoH models, greatly improving the applicability of the proposed MoH method. In contrast, MoA directly combines multi-head attention with MoE. Due to the adoption of shared keys and values, MoA must be trained from scratch, which limits its applicability. Finally, in terms of model frameworks, our MoH is validated across various popular model frameworks and tasks, including ViT, DiT, and decoder-only LLMs, while MoA is only validated for language tasks.
@@ 6 Conclusion
[…]
@@ Acknowledgements
Abstract.   This appendix provides additional discussions (Appendix A), implementation details (Appendix B), several additional experiments (Appendix C), more qualitative analysis (Appendix D), and details of quantitative evaluations for LLMs (Appendix E).
@@ A.1 Why is MoH Superior to Vanilla Multi-Head Attention?
In contrast, in MoH, routed heads are trained only on smaller subsets of data specifically assigned to them. In MoH’s routing mechanism, the data is divided into $h-h_{s}$ subsets $\{D_{1},D_{2},...,D_{h-h_{s}}\}$, with each subset corresponding to a routed head. Besides, the routing score for each attention head acts as an adaptive adjustment to the learning rate, enabling the attention heads in MoH to specialize more effectively. Given a minibatch of data $D$ and the router G()G(*), the gradient of each routed head in MoH can be written as $\mathbb{E}_{x\in D_{i}}[G(x)_{i}\frac{\partial\mathcal{L}(x)}{\partial h_{i}}]$. The gradient of each shared head in MoH can be written as $\mathbb{E}_{x\in D}[G(x)_{i}\frac{\partial\mathcal{L}(x)}{\partial h_{i}}]$. As shown in Tab. A, the routing mechanism and adaptive weights in MoH enable attention heads to specialize more effectively compared to standard multi-head attention.
@@
[…]
@@ A.2 Limitations and Future Work
[…]
Multimodal Inputs.   Effectively processing information from multiple modalities in the attention mechanism remains an open question. Recent work (wan2024look) has shown that visual and textual tokens exhibit distinct attention patterns in multi-head attention. Future work could explore the attention patterns of MoH with different modal inputs, for example within multimodal large language models (jin2024chat; lin2023video; lin2024moe; liusphinx; jin2023video; jin2024hierarchical).
@@ B.1 ViT for Image Classification
Training Details.   Our MoH-ViT models are trained for 300 epochs using automatic mixed precision across 8 GPUs. We follow the training strategy of TransNeXt, which includes various data augmentation techniques, including Random Augmentation (cubuk2020randaugment), Mixup (zhang2017mixup), CutMix (yun2019cutmix), and Random Erasing (zhong2020random). We also apply Label Smoothing (szegedy2016rethinking) and DropPath (huang2016deep) to regularize our models. We optimize our models using AdamW optimizer (loshchilov2017decoupled) with a gradient clipping norm of 1.0 and a weight decay of 0.05. The initial learning rate is set to 1e-3, with a 5-epoch warm-up starting at 1e-6. A cosine learning rate scheduler (loshchilov2016sgdr) is employed to decay the learning rate. During training, images are randomly cropped to a size of $\times224$. It is worth noting that we do not use Exponential Moving Average (EMA) weights.
@@ B.2 DiT for Class-Conditional Image Generation
Training Details.   Following DiT, the final linear layer is initialized with zeros, and all other layers follow standard ViT weight initialization. We train all models using the AdamW optimizer (loshchilov2017decoupled) with a constant learning rate of 1e-4, no weight decay, and a batch size of 256, applying horizontal flips for data augmentation. Following DiT, we employ the Exponential Moving Average (EMA) of MoH-DiT weights during training with a decay rate of 0.9999, generating all images using the EMA model. We use an off-the-shelf pre-trained variational autoencoder (kingma2013auto) model from Stable Diffusion (rombach2022high). Following TransNeXt, our attention-head activation budget is unevenly distributed across layers, with fewer attention heads activated in the shallow layers and more in the deeper layers.
@@ B.3 Training LLMs from Scratch
Model Settings.   For training LLMs from scratch, we use Megatron (shoeybi2019megatron), an open-source training code, as the training framework. The detailed hyper-parameter settings of various MoH-LLMs are shown in Tab. C.
@@
  • The RedPajama (together2023redpajama) includes training data from seven domains: CommonCrawl, C4, Github, Wikipedia, Books, ArXiv, and StackExchange.
  • The Dolma (dolma), a large and diverse open English text corpus, contains 3 trillion tokens sampled from seven sources, including web pages from Common Crawl, code from The Stack, curated web data from C4 (raffel2020exploring), social media conversations from Reddit, academic papers from PeS2o, public domain books from Project Gutenberg, and comprehensive content from Wikipedia and Wikibooks.
  • The Pile (pile), an open-source English text corpus for training large language models, includes 22 diverse, publicly available datasets such as Wikipedia, NIH ExPorter, ArXiv, Books3, BookCorpus2, OpenSubtitles, YoutubeSubtitles, and Enron Emails.
@@
Training Hyper-Parameters.   Tab. E shows the detailed training hyper-parameters of MoH-LLMs. Specifically, all MoH-LLMs are trained with the AdamW optimizer (loshchilov2017decoupled), using a batch size of 4 million tokens with a sequence length of 2048. The final learning rate is set to 10% of the maximum. During training, a weight decay of 0.1 and gradient clipping of 1.0 are applied. For LLM-S and MoH-LLM-S, the maximum learning rate is set to 3e-4. For LLM-B and MoH-LLM-B, the maximum learning rate is set to 5e-4.